# Variable Elliptic Structures on the Plane: Transport Dynamics, Rigidity, and Function Theory

### v3

@@ 9.5.8 Remarks
Canonical metric (flatness)
The canonical Riemannian metric associated with the principal symbol is written in terms of $\lambda=p+iq$ as
$ds^{2}=\frac{1}{\Im\lambda}\Bigl(dx^{2}+2\,\Re\lambda\,dx\,dy+|\lambda|^{2}\,dy^{2}\Bigr).$
This metric depends only on $(\alpha,\beta)$ and coincides with the metric induced by the rigid elliptic algebra. Let $K$ denote its Gaussian curvature.
The principal symbol is called flat if
$K\equiv 0.$
Summary of intrinsic invariants
With this alignment, the two intrinsic invariants of a real first–order elliptic system are:
  • the transport invariant
    $T=\lambda_{x}+\lambda\,\lambda_{y},$
    measuring rigidity;
  • the Gaussian curvature $K$ of the canonical metric, measuring flatness.
Both invariants depend only on the principal symbol and are invariant under changes of dependent variables and orientation–preserving coordinate changes.
All subsequent statements, computations, and proofs in this chapter are understood with $\lambda$ as the spectral slope associated to the elliptic structure $(\alpha,\beta)$, in full consistency with the Vekua formulation.
10.1 The complex root and intrinsic field
The equation $Q(1,\tau)=0$ has two complex conjugate roots. We select the one with positive imaginary part:
$\tau=p+iq=\frac{-\beta+i\sqrt{4\alpha-\beta^{2}}}{2},\qquad q>0.$
This complex-valued function $\tau(x,y)$ encodes the intrinsic geometry of the principal symbol.
10.1.1 Two intrinsic invariants
Two fundamental invariants arise naturally from $\tau$.
1. The canonical metric.
Define
$ds^{2}=\frac{1}{q}\Bigl(dx^{2}+2p\,dx\,dy+(p^{2}+q^{2})\,dy^{2}\Bigr).$
This Riemannian metric depends only on the principal symbol. Let $K$ denote its Gaussian curvature.
2. The transport invariant.
Define the first-order quantity
$T=\tau_{x}+\tau\tau_{y}.$
This expression measures how the symbol varies along its own characteristic direction. It depends only on the first derivatives of $\tau$ and hence only on the principal symbol.
10.1.2 Rigidity and flatness
We introduce two geometric conditions on the principal symbol:
  • The system is rigid if $T\equiv 0$.
  • The system is flat if the canonical metric has $K\equiv 0$.
Both properties are intrinsic and independent of lower-order terms of the PDE.
10.2 Statement of the theorem
Theorem 10.1 (Rigidity–Flatness Theorem).
Let $\mathcal{L}$ be a smooth real first–order elliptic system on a connected planar domain $\Omega$. If its principal symbol is rigid ($T=0$) and flat ($K=0$), then the complex root $\tau$ is constant on $\Omega$. Equivalently, the coefficients $\alpha$ and $\beta$ of the structure polynomial are constant, so the principal symbol is locally equivalent to that of a constant–coefficient elliptic system.
10.2.1 Jet formulation
Write $\tau=p+iq$ with $q>0$. At a point $z\in\Omega$, the 22–jet of $(p,q)$ is
$j^{2}_{z}(p,q)=(p,q,p_{x},p_{y},q_{x},q_{y},p_{xx},p_{xy},p_{yy},q_{xx},q_{xy},q_{yy})|_{z}.$
All second–order invariants of the symbol, including $K$, are rational functions of this jet.
Rigidity $T=0$ is equivalent to
$p_{x}=-p\,p_{y}+q\,q_{y},\qquad q_{x}=-p\,q_{y}-q\,p_{y}.$ (10.1)
Differentiating (10.1) eliminates all xx–derivatives of order $\leq 2$. On the rigid locus, the free jet variables reduce to
$(p,q,p_{y},q_{y},p_{yy},q_{yy}).$
10.2.2 Jet-level curvature identity
Define the curvature numerator
$N:=-2q^{3}K.$
Lemma 10.2 (Rigid curvature identity).
On the rigid locus,
$N=-q\,\Lambda(p,q)\,q_{yy}+R(p,q;p_{y},q_{y}),$
where $\Lambda(p,q)>0$ for $q>0$, and
$R=A(p,q)p_{y}^{2}+2B(p,q)p_{y}q_{y}+C(p,q)q_{y}^{2}$
is positive definite.
After clearing denominators this becomes a polynomial identity in finitely many jet variables and can therefore be symbolically verified.
10.2.3 Consequences of flatness
If $T=0$ and $K=0$, then $N=0$, so
$q\,\Lambda(p,q)\,q_{yy}=R(p,q;p_{y},q_{y})\geq 0.$
10.2.4 Characteristic representation
Rigidity implies the complex Burgers equation
$\tau_{x}+\tau\tau_{y}=0.$
Introduce characteristic coordinates
$\zeta=y-\tau x.$
Locally there exists $C^{1}$ $\Phi$ such that
$\tau(x,y)=\Phi(\zeta).$
Differentiating,
$\tau_{y}=\frac{\Phi^{\prime}}{1+x\Phi^{\prime}},\qquad\tau_{yy}=\frac{\Phi^{\prime\prime}}{(1+x\Phi^{\prime})^{3}}.$
10.2.5 Degree argument
Restrict to a fixed characteristic slice $\zeta=\mathrm{const}$. Then $\Phi^{\prime}$ and $\Phi^{\prime\prime}$ are constants with respect to $x$. Multiplying $N=0$ by $|1+x\Phi^{\prime}|^{6}$ produces a polynomial in $x$. The term involving $q_{yy}$ has degree at most $3$, while the quadratic form $R$ produces degree $4$. Since $R$ is positive definite, the leading coefficient is proportional to $|\Phi^{\prime}|^{6}$ and must vanish. Hence $\Phi^{\prime}=0$.
Therefore $\tau_{y}\equiv 0$, and rigidity gives $\tau_{x}=0$. Thus $\tau$ is constant.
Proof of Theorem 10.1.
Combine Lemma 10.2 with the characteristic representation and the degree argument above. ∎
Conclusion
This monograph set out from a simple but decisive shift in viewpoint: the imaginary unit should not be treated as a fixed algebraic constant, but as intrinsic geometric data. Allowing the generator $i(x,y)$ of a rank–two real algebra bundle to vary forces the algebra itself to become part of the geometry of the plane. Once this step is taken, differentiation of the structure is no longer optional, and the derivatives $i_{x}$ and $i_{y}$ emerge as genuine geometric quantities rather than formal artifacts.
The central object organizing the theory is the intrinsic obstruction
$G:=i_{x}+i\,i_{y}.$
This quantity appears prior to any analytic notion of holomorphicity. It measures the incompatibility between differentiation in the base variables and multiplication in the fibers, and it governs, simultaneously and unavoidably, all deviations from the classical theory: the inhomogeneity of generalized Cauchy–Riemann systems, the failure of the Leibniz rule, and the appearance of source terms in first–order structure equations. These phenomena are not imposed by analytic assumptions; they are forced by the geometry of a moving elliptic algebra.
A first structural outcome of the theory is that variability of the coefficients $(\alpha,\beta)$ is not arbitrary. Eliminating the derivatives of the generator from the structure reduction leads to a universal transport law for the spectral parameter
@@ Conclusion
which takes the form of a (possibly forced) complex inviscid Burgers equation written in the standard complex algebra. This transport equation belongs to the structural layer of the theory: it arises solely from allowing the elliptic structure to vary and does not rely on integrability, coordinate trivialization, or function–theoretic input.
Within this universal transport framework, rigidity occupies a precise and limited role. The condition
$i_{x}+i\,i_{y}=0$
does not generate the dynamics; it selects the conservative regime of the already–existing transport law. In this regime, the Burgers equation becomes unforced, and several independent obstructions collapse simultaneously: the generalized Cauchy–Riemann operator becomes a derivation, the real first–order system becomes homogeneous, and the generator is transported compatibly with its own multiplication law. Rigidity is therefore a compatibility condition, not a simplifying assumption.
The explicit $\varepsilon$–family constructed in this work demonstrates that rigidity does not force triviality. Genuinely nonconstant elliptic structures exist, although ellipticity naturally restricts the domain of definition. From the transport viewpoint, these restrictions reflect characteristic phenomena familiar from first–order hyperbolic systems and are not pathologies of the theory.
On the analytic side, rigidity is exactly the threshold at which a coherent function theory re–emerges. In the rigid regime, a Cauchy–Pompeiu formula adapted to variable elliptic structures persists. The natural Cauchy 11–form is not closed, and its non–closure produces an unavoidable interior correction term. Rather than being an artifact, this correction identifies a canonical covariant Cauchy–Riemann operator whose null space provides the appropriate notion of holomorphicity in the variable setting. Although covariantly holomorphic sections are not closed under pointwise multiplication, a scalar gauge transformation restores closure through a natural weighted product. Power series representations follow from this framework, but only after freezing the algebra at the expansion point, making explicit where classical arguments genuinely fail in the variable case.
Finally, the theory interfaces naturally with the classical Vekua program. General first–order elliptic systems induce variable elliptic structures, and their reduction to Vekua form reveals the intrinsic obstruction as the true geometric source of inhomogeneity. Rigidization, formulated as an ordinary differential equation along characteristic curves, provides a constructive mechanism to pass from non–rigid to rigid structures without solving global Beltrami equations. In the rigid regime, sharp similarity principles, initial value problems, and algorithmic solution schemes become available, even in the presence of non–uniform ellipticity.
The scope of this work has been deliberately local and explicit. Rather than appealing to abstract integrability frameworks, the theory has been developed at the level of concrete first–order computations in order to keep the geometric and dynamical mechanisms fully visible. From this perspective, variable elliptic structures exhibit a clear hierarchy: transport is fundamental, rigidity is selective, and analytic closure is a distinguished consequence rather than an axiom.
This hierarchy reframes classical complex analysis as a special case of a broader geometric theory.
@@
Here, the algebra begins to move.
@@ E.5 Conclusion
Appendix F Symbolic Verification of the Rigid Second–Order Expansion

### v4

@@ Conclusion
is precisely the horizontal component of the pullback of the Maurer–Cartan form of $S^{3}$. No new invariant has appeared. The obstruction identified through algebraic elimination is the part of the conjugation derivative that cannot be absorbed by a vertical gauge change.
The variability of the coefficients $(\alpha,\beta)$ is governed by a universal transport law. Eliminating $i_{x}$ and $i_{y}$ from the differentiated quadratic relation forces the spectral parameter
@@
to satisfy a (possibly forced) complex inviscid Burgers equation in the standard complex algebra. Rigidity,
$G=0,$
selects the regime in which the horizontal component vanishes. In that regime, the Burgers equation becomes conservative, the generalized Cauchy–Riemann operator becomes a derivation, and the generator is transported compatibly with its own multiplication. Rigidity is the condition that the motion of the imaginary unit is purely vertical in the Hopf bundle.
On the analytic side, rigidity marks the threshold at which a coherent function theory reappears. A Cauchy–Pompeiu formula persists, a covariant Cauchy–Riemann operator emerges canonically, weighted products restore multiplicative closure, and power series representations follow after freezing the algebra at a point. The variable theory does not abolish complex analysis; it reveals which of its properties depend on rigidity and which depend on the hidden constancy of the standard structure.
The translation to the Beltrami variable
$\mu=\frac{\lambda-i}{\lambda+i}\in\mathbb{D}$
recasts the entire theory in the language of the Poincaré disk. The rigidity condition $G=0$ becomes the self–dilatation equation
$\mu_{\bar{z}}+\mu\,\mu_{z}=0,$
which states that $\mu$ is a dilatational map with respect to its own structure. The Poincaré metric, which appeared uninvited in the convergence theory as the normalization $b_{\min}^{-2}$, turns out to be the intrinsic geometry of the parameter space $\mathbb{D}$.
From this vantage, the central discovery of the monograph becomes visible. The difficulty of a variable elliptic structure lives in a plane, not on a line. Two invariants govern it: the Beltrami modulus $\|\mu\|_{C^{0}}$, a zeroth–order quantity measuring where the structure sits in the Poincaré disk, and the Poincaré residual $R(\mu)=\mu_{\bar{z}}+\mu\,\mu_{z}$, a first–order quantity measuring how the structure moves through it. These are independently prescribable. All four quadrants of the difficulty plane $\bigl(\|\mu\|_{C^{0}},\,\|R\|_{C^{0,\alpha}}\bigr)$ are populated by smooth elliptic structures.
The classical Beltrami literature—the Measurable Riemann Mapping Theorem, the Ahlfors–Bers theorem, the theory of quasiconformal mappings—projects this plane onto the horizontal axis. Only $\|\mu\|_{C^{0}}$ survives. The vertical axis, the transport field, is invisible. A rigid structure at $|\mu|=0.99$ supports a complete function theory. A non–rigid structure at $|\mu|=0.01$ does not, without first being rigidized. The zeroth–order invariant cannot see the first–order distinction.
The explicit families constructed in this work make this concrete. The $\varepsilon$–family demonstrates that rigidity does not force constancy: nontrivial elliptic structures correspond to genuine motions on $S^{2}$, though ellipticity naturally restricts the domain through characteristic phenomena. The $\delta$–family demonstrates that zeroth–order difficulty and first–order difficulty are decoupled: $\|\mu\|_{C^{0}}\to 1$ while $R\equiv 0$ identically. A solver that measures difficulty by $|\mu|$ alone sees a problem approaching maximal degeneracy. The rigidizer sees a problem already solved.
The p(x)p(x)–analytic theory, developed by the Bers school and its successors, populates the real diameter of the Poincaré disk: $\mu=(\sqrt{p}-1)/(\sqrt{p}+1)\in(-1,1)$. Along this diameter, $\beta=0$ identically; the variable exponent modulates conformal scaling but never opens the transport channel. The relationship is exact:
$\begin{array}[]{ccccc}\text{real analysis}&\xrightarrow{\quad\text{analytic continuation}\quad}&\text{complex analysis}\\[6.0pt] p(x)\text{-analytics}&\xrightarrow{\quad\beta\neq 0\quad}&\text{VES}\end{array}$
The generating pairs, successor structures, embedding theorems, and regularity results of the established theory form the boundary data from which the full variable theory departs. The observation is not that these results are insufficient. It is that they explore the real diameter of a disk whose off–axis structure was never investigated.
The residual budget $B(\mu)$ at each point of the disk measures how far the constructive theory reaches: the maximum Poincaré residual for which the rigidization scheme converges. The budget is not rotationally symmetric. It reflects an asymmetry intrinsic to the Möbius transformation relating the spectral parameter to the Beltrami variable. The constructive envelope in the difficulty plane is therefore a computable region, not a universal constant.
The diagnostic algorithm that emerges from this analysis classifies first–order elliptic systems into three branches: structures within the constructive envelope, accessible to the full VES toolkit; p(x)p(x)–analytic structures on the real diameter, governed by the Bers theory; and structures with genuine transport obstruction beyond the envelope, requiring classical quasiconformal methods. Without the independence of position and velocity in the Poincaré disk, this triage is impossible.
Throughout, the development has remained local and first–order. The geometry of the Hopf bundle was not required to derive the theory; it confirms, at the end, what the algebra had already forced. The Poincaré disk was not assumed as a framework; it emerged from the convergence theory as the natural geometry of the parameter space. The independence of the two axes of difficulty was not imposed by construction; it follows from the elementary fact that a function and its derivative are independent objects.
Transport is fundamental. Rigidity is selective. Analytic closure is consequential.
@@
In Hopf coordinates we see what was moving: a rotating imaginary unit on $S^{2}$.
In Beltrami coordinates we see what was hidden: a two–dimensional difficulty space projected onto a line.
Here, the algebra moves.
@@ E.5 Conclusion
Appendix F Quaternionic Reformulation of Rigidity via the Hopf Fibration
F.1 Motivation
A variable elliptic structure on a domain $\Omega\subset\mathbb{R}^{2}$ is encoded by a generator $i=i(x,y)$ satisfying the structure reduction
$i^{2}+\beta(x,y)\,i+\alpha(x,y)=0,$ (F.1)
with ellipticity discriminant
$\Delta(x,y):=4\alpha(x,y)-\beta(x,y)^{2}>0.$ (F.2)
The intrinsic obstruction is
$G:=i_{x}+i\,i_{y},$ (F.3)
and the structure is called rigid precisely when $G\equiv 0$.
The purpose of this appendix is to recast rigidity in a purely quaternionic language via the Hopf fibration, clarifying the geometric content of $G$.
F.1.1 Normalization to a unit imaginary quaternion
Completing the square in (F.1) yields
$(2i+\beta)^{2}=\beta^{2}-4\alpha=-\Delta.$
Define the normalized generator
$u(x,y):=\frac{2i(x,y)+\beta(x,y)}{\sqrt{\Delta(x,y)}}.$ (F.4)
Then
$u^{2}=-1,$ (F.5)
so $u$ takes values in the unit imaginary sphere
$S^{2}=\{\,q\in\operatorname{Im}\mathbb{H}:\ |q|=1\,\}\subset\mathbb{H}.$
Thus, an elliptic quadratic structure canonically determines a smooth map $u:\Omega\to S^{2}$.
F.2 Hopf lift and gauge freedom
Fix a constant imaginary unit $i_{0}\in\operatorname{Im}\mathbb{H}$ with $i_{0}^{2}=-1$. Since every $u(x,y)\in S^{2}$ is conjugate to $i_{0}$, locally there exists a Hopf lift $q:\Omega\to S^{3}$ (unit quaternions) such that
$u=q\,i_{0}\,q^{-1}.$ (F.6)
The lift is unique up to the $S^{1}$-gauge transformation
$q\ \mapsto\ q\,e^{i_{0}\theta(x,y)}.$ (F.7)
F.3 Maurer–Cartan form and differentiation
Define the left-trivialized Maurer–Cartan components
$\Omega_{x}:=q_{x}q^{-1},\qquad\Omega_{y}:=q_{y}q^{-1},$ (F.8)
which lie in $\mathfrak{su}(2)\cong\operatorname{Im}\mathbb{H}$. Differentiating (F.6) and using $(q^{-1})_{x}=-q^{-1}q_{x}q^{-1}$ gives
$u_{x}=[\Omega_{x},u],\qquad u_{y}=[\Omega_{y},u],$ (F.9)
where [A,B]=ABBA[A,B]=AB-BA denotes the commutator in $\mathbb{H}$.
F.3.1 Vertical–horizontal decomposition
At each point, $\mathfrak{su}(2)\cong\operatorname{Im}\mathbb{H}$ splits as
$\operatorname{Im}\mathbb{H}=\mathrm{span}\{u\}\ \oplus\ u^{\perp}.$
Accordingly decompose
$\Omega_{x}=\Omega_{x}^{\mathrm{ver}}+\Omega_{x}^{\mathrm{hor}},\qquad\Omega_{y}=\Omega_{y}^{\mathrm{ver}}+\Omega_{y}^{\mathrm{hor}},$ (F.10)
where $\Omega^{\mathrm{ver}}\parallel u$ and $\Omega^{\mathrm{hor}}\perp u$. Since $\Omega^{\mathrm{ver}}$ commutes with $u$, it follows from (F.9) that
$u_{x}=[\Omega_{x}^{\mathrm{hor}},u],\qquad u_{y}=[\Omega_{y}^{\mathrm{hor}},u].$ (F.11)
Under the gauge change (F.7), the Maurer–Cartan form changes by adding a vertical term, so the horizontal components and the commutators in (F.11) capture the gauge-invariant variation of $u$.
F.4 Quaternionic form of the obstruction
The obstruction $G=i_{x}+i\,i_{y}$ can be rewritten in terms of $u$. Up to the scalar normalizations in (F.4), the intrinsic combination governing deviations from the constant theory corresponds to
$\mathcal{G}:=u_{x}+u\,u_{y}.$ (F.12)
Using (F.11), we may express $\mathcal{G}$ in Hopf variables as
$\mathcal{G}=[\Omega_{x}^{\mathrm{hor}},u]\;+\;u\,[\Omega_{y}^{\mathrm{hor}},u].$ (F.13)
This expression is gauge-invariant (it is insensitive to vertical modifications of the lift).
F.5 Quaternionic rigidity theorem
Theorem F.1 (Quaternionic rigidity).
Let $u:\Omega\to S^{2}\subset\operatorname{Im}\mathbb{H}$ be the normalized generator (F.4) associated with an elliptic quadratic structure (F.1). Let $q$ be a local Hopf lift (F.6) with Maurer–Cartan components (F.8) and decomposition (F.10). Then the following are equivalent:
  1. (i)
[…]
    $u_{x}+u\,u_{y}=0\quad\text{on }\Omega.$ (F.14)
  2. (ii)
    (Horizontal Maurer–Cartan cancellation)
[…]
  3. (iii)
    (Purely vertical transport along the characteristic direction) the pulled-back Maurer–Cartan form has no effective horizontal component along the direction $\partial_{x}+u\,\partial_{y}$, i.e. $u$ undergoes only fiber (Hopf) gauge rotation along that direction.
Moreover, (F.14) is the Hopf/quaternionic avatar of the original rigidity condition $G=i_{x}+i\,i_{y}=0$.
Proof.
By (F.11), we have $u_{x}=[\Omega_{x}^{\mathrm{hor}},u]$ and $u_{y}=[\Omega_{y}^{\mathrm{hor}},u]$. Substituting these into (F.12) yields (F.13), and hence $\mathcal{G}\equiv 0$ is equivalent to (F.15). The gauge statement follows from the fact that changing the lift by (F.7) alters only the vertical part of the Maurer–Cartan form, leaving the horizontal commutator expressions unchanged. ∎
F.6 Interpretation
Equation (F.14) says that the $S^{2}$-valued imaginary direction field $u$ is transported compatibly with its own multiplication rule: along the characteristic direction $\partial_{x}+u\,\partial_{y}$, the field exhibits no genuine horizontal twisting in $S^{2}$, only Hopf fiber rotation. In this sense rigidity is a horizontality cancellation condition in the Hopf fibration rather than an analytic assumption.
Appendix G Rigidization as a Computational Normal Form
Purpose of this Appendix
This appendix clarifies the intended scope of rigidization within the theory of variable elliptic structures. Rigidization is not proposed as a replacement for global uniformization (Beltrami theory), nor as a universal flattening mechanism. Instead, it is designed as a computationally optimized normal form that exploits intrinsic transport structure when present, and that can be used either as a standalone reduction or as a preprocessing step for uniformization-based solvers.
The guiding philosophy is:
Rigidization removes exactly the analytic obstructions that prevent the use of the constant-coefficient toolbox, while preserving geometric variability that is irrelevant for computation but expensive to eliminate.
G.1 Rigidization versus Uniformization
Let $\lambda(x,y)$ denote the spectral slope of a variable elliptic structure, with intrinsic obstruction
$T:=\lambda_{x}+\lambda\,\lambda_{y}.$
Uniformization (Beltrami)
Uniformization seeks a diffeomorphism $\Phi$ such that the pullback structure is constant. This solves the maximal geometric problem but requires solving a globally coupled nonlinear elliptic PDE. From a computational standpoint, uniformization is:
  • global and nonlocal,
  • expensive to compute,
  • insensitive to existing transport structure,
  • often unnecessary when only analytic compatibility is required.
Rigidization
Rigidization seeks a diffeomorphism $\Phi$ such that the pullback structure satisfies
$\widetilde{T}=0,\qquad\text{i.e.}\quad\widetilde{\lambda}_{X}+\widetilde{\lambda}\,\widetilde{\lambda}_{Y}=0,$
but does not require $\widetilde{\lambda}$ to be constant.
This is a strictly weaker requirement than uniformization. Nevertheless, once rigidity is achieved:
  • the Leibniz defect vanishes,
  • the generalized Cauchy–Riemann system becomes homogeneous,
  • Cauchy–Pompeiu and similarity principles apply,
  • the full constant-structure analytic toolkit becomes available.
Thus rigidization reaches the minimal normal form required for analysis, without over-solving the geometric problem.
G.2 Why “Weaker” Means “Easier”
The rigidization condition is first-order in the structure and second-order in the coordinates, and its linearization is an elliptic system whose principal symbol is explicitly computable. As shown in Chapter 10, this enables:
  • local Newton convergence under a quantitative small-obstruction condition (Theorem 10.15),
  • inner solvers based on transport–wave splitting (shifted ADI with Wachspress parameters),
  • direct control of analytic compatibility without global flattening.
In contrast, full uniformization removes curvature and transport simultaneously, even when curvature plays no analytic role for the PDE being solved.
From a solver perspective, rigidization is therefore cheaper, more local, and better aligned with problems that already possess a preferred transport geometry.
G.3 A Solver-Oriented Triage
Rigidization is intended to be used selectively. Three intrinsic indicators guide solver choice:
  1. 1.
    Obstruction magnitude:
    $\rho_{T}:=\frac{\|T\|_{C^{0,\alpha}}}{b_{\min}^{2}},\qquad b_{\min}:=\inf\operatorname{Im}\lambda.$
  2. 2.
    Metric curvature magnitude: a scale-normalized curvature indicator $\rho_{K}$ for the canonical metric.
  3. 3.
    Transport dominance: a problem-dependent measure of whether a preferred first-order flow direction is present.
These lead to the following computational policy:
  • Small obstruction ($\rho_{T}\ll 1$): Rigidize immediately using the Newton scheme of Chapter 10. This is the regime covered by Theorem 10.15 with full convergence guarantees.
  • Moderate obstruction, transport-dominated PDE: Attempt rigidization via a continuity strategy (interpolating from a rigid reference structure $\lambda_{0}=i$ via $\lambda^{t}=(1-t)\lambda_{0}+t\lambda$, taking incremental steps in $t$ with each step small enough for the Newton iteration of Theorem 10.15 to converge). If the obstruction is substantially reduced, proceed in rigid coordinates; otherwise fall back to uniformization.
  • Large obstruction, negligible curvature: Use Beltrami uniformization directly.
  • Diffusion-dominated problems: Use uniformization; rigidization offers no advantage when no transport structure is present.
This policy is not ad hoc: it reflects the analytic role played by the obstruction in the failure of derivation properties, even though the specific thresholds and decision boundaries are solver-dependent.
G.4 Rigidization as a Preconditioner
Even when rigidization does not succeed globally, it can still be valuable.
Chapter 10 shows that naive transport–wave splitting does not converge as a standalone iteration, but is effective as an inner solver for the linearized elliptic problem. This mirrors standard numerical practice:
  • rigidization removes transport incompatibility first,
  • uniformization then operates on a better-conditioned structure,
  • overall solver stability and convergence improve.
In this sense, rigidization acts as a geometric preconditioner for uniformization-based methods.
More concretely, the incremental continuity strategy described above provides a practical algorithm even without a closed-form global existence theorem: one advances in $t$ until the Newton step fails to converge, records the partially rigidized structure, and hands the residual obstruction to a uniformization solver operating on a problem whose effective $\rho_{T}$ has been reduced.
G.5 The Global Rigidization Problem
It is natural to ask whether rigidization can always be achieved on compact domains, without any smallness condition on the obstruction $T$.
Conjecture 1 (Global rigidization).
Let $\Omega\subset\mathbb{R}^{2}$ be a compact simply connected domain with $C^{2,\alpha}$ boundary. Let $\lambda\in C^{2,\alpha}(\bar{\Omega})$ with $\operatorname{Im}\lambda\geq b_{\min}>0$ on $\bar{\Omega}$. Then there exists a $C^{2,\alpha}$ diffeomorphism $\Phi:\bar{\Omega}\to\Phi(\bar{\Omega})$ such that the pullback structure is rigid.
A natural proof strategy is the continuity method: interpolating from a rigid reference $\lambda_{0}=i$ to the target $\lambda$, the success set is nonempty (at $t=0$) and open (by Theorem 10.15). Closedness would follow from uniform a priori estimates on the rigidizing diffeomorphisms—specifically, a uniform $C^{2,\alpha}$ bound and a uniform Jacobian lower bound along the family.
These a priori estimates constitute the core difficulty. Three coupled sub-problems are involved:
  1. 1.
    Jacobian lower bound. Show that $\inf_{\bar{\Omega}}J_{\Phi}\geq c>0$ for any rigidizing diffeomorphism, with $c$ depending only on $b_{\min}$, $\|\lambda\|_{C^{2,\alpha}}$, and $\Omega$. A promising route is to derive a second-order elliptic equation (or inequality) for $\log J_{\Phi}$ from the ratio condition and apply the maximum principle, using the boundary normalization $\Phi|_{\partial\Omega}=\mathrm{id}$ (so $J=1$ on $\partial\Omega$).
  2. 2.
    Burgers constraint on compact domains. The pullback $\widetilde{\lambda}$ solves conservative Burgers, whose solutions are constant along characteristics. On compact domains with $\operatorname{Im}\widetilde{\lambda}>0$, the characteristic map from boundary data to the interior is expected to be injective (no shock formation in the elliptic regime). A correct proof requires a quantitative injectivity estimate on the characteristic map—specifically, a determinant lower bound on the map $(s,X)\mapsto(X,Y)$ involving $\partial\widetilde{\lambda}/\partial s$ and domain geometry, not merely the condition $\operatorname{Im}\widetilde{\lambda}>0$.
  3. 3.
    $C^{2,\alpha}$ a priori bound. With the Jacobian controlled, the rigidization system is uniformly elliptic. Schauder estimates (Agmon–Douglis–Nirenberg) should yield global $C^{2,\alpha}$ bounds, but the application to this specific quasilinear $2\times 2$ system requires verifying the structural conditions under which elliptic regularity and boundary estimates hold.
These sub-problems are mutually dependent: the Burgers constraint feeds into the Jacobian bound, which in turn feeds into the Schauder estimates. As of this writing, none of the three has been rigorously closed, and the conjecture remains open.
We record it here both as an invitation to further work and as context for the solver-oriented approach of this appendix: the computational triage and preconditioner strategies above are designed to be effective regardless of whether Conjecture 1 is eventually proved.
G.6 What Is Not Claimed
To be explicit about the boundaries of the current theory, this monograph does not claim:
  • global rigidization for arbitrary elliptic structures on compact domains (this remains an open problem; see Conjecture 1),
  • a priori Jacobian lower bounds independent of the obstruction magnitude (the estimates in Chapter 10 depend on the smallness condition $\|T\|<C\,b_{\min}^{2}$),
  • replacement of Beltrami theory as a universal tool.
These are not limitations of design intent but reflections of the current state of the analysis. The global theory may well be within reach—the structural ingredients (ellipticity of the linearized problem, Burgers characteristic geometry, similarity-principle arguments) are all present—but assembling them into a complete a priori package requires new ideas, particularly for the control of the Jacobian.
G.7 Summary
Rigidization should be viewed as:
  • a maximal computable normal form,
  • a solver-side exploitation of transport geometry,
  • a selective alternative or complement to uniformization,
  • a mechanism that preserves just enough geometry to be efficient.
When it applies, rigidization delivers the full analytic power of the constant-coefficient theory at a fraction of the computational cost. When its applicability has not been established, it still provides valuable structural diagnostics and preconditioning for heavier methods.
In this sense, rigidization is not a competitor to uniformization but a practical intermediary between raw variable geometry and full geometric flattening—one whose global reach remains a compelling open question in its own right.
Appendix H Symbolic Verification of the Rigid Second–Order Expansion